The ABC corporation pension plan provides a lifetime annual income to its employees upon retirement at age 65. The plan provides 3% for each year of service of the average of the employee's earnings for their last 2 years on the job. Mark retires at age 65 with 16 years of service. His earnings for his last 2 years were $55465 and $57760. What is his annual pension benefit? Round your answer to the nearest dollar.
Answer #1

average of earning of last 2 years=(55465+57760)/2=56612.5

amount of annual pension benefit =56612.5*3%*16 years= $27714

monthly installment=27714/12=2264.5 or 2265(rounded off)